OpenStudy (anonymous):

Graph the first six terms of a sequence where a1 = -10 and d = 3.

OpenStudy (solomonzelman):

graph the terms in a sequence? Well your x and y coordinates in general would look as follows: \(\large\color{black}{ \displaystyle (1,~a_1) }\) \(\large\color{black}{ \displaystyle (2,~a_2) }\) \(\large\color{black}{ \displaystyle (3,~a_3) }\) \(\large\color{black}{ \displaystyle (4,~a_4) }\) and on...

OpenStudy (solomonzelman):

with d=3, wat would be your second term ?

OpenStudy (anonymous):

where do you put the d?

OpenStudy (solomonzelman):

the "d" is a common difference. Can I give you an example of a different sequence >?

OpenStudy (anonymous):

yeah whatever helps please!

OpenStudy (solomonzelman):

Find the first four terms of a sequence where \(\large\color{black}{ \displaystyle a_1= 4 }\) and \(\large\color{black}{ \displaystyle d=2 }\). your first term (denoted as \(\large\color{black}{ \displaystyle a_1 }\) ) is 4. your second term (denoted as \(\large\color{black}{ \displaystyle a_2 }\) ) is: \(\large\color{black}{ \displaystyle a_2=a_1+d }\) you know \(\large\color{black}{ \displaystyle a_1=4 }\) and \(\large\color{black}{ \displaystyle d=2 }\) \(\large\color{black}{ \displaystyle a_2=4+2=6 }\)

OpenStudy (solomonzelman):

So, to find each preceding term you will be adding this "d" (in this case 2).

OpenStudy (solomonzelman):

So the sequence in general is: \(\large\color{black}{ \displaystyle a_1,~~~(a1)+d,~~~(a1+d)+d,~~~(a1+d+d)+d, }\) ... etc

OpenStudy (solomonzelman):

ok, lets (I guess) do your example to save some time.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_1=-10}\) with d=3

OpenStudy (anonymous):

so it'd be -10+3?

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ \displaystyle a_2=-10+3=7}\)

OpenStudy (solomonzelman):

sorry not 7, but -7

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_2=-10+3=-7}\)

OpenStudy (solomonzelman):

then \(\large\color{black}{ \displaystyle a_3=-7+3=-4}\)

OpenStudy (solomonzelman):

can you tell me the \(\large\color{black}{ \displaystyle a_4}\) ?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

so if your starting with a2, does that mean to make them points on a grap it'd be (1,-10)(2,-7)(3,-4) etc...??

OpenStudy (solomonzelman):

yes \(\large\color{black}{ \displaystyle a_4=-1}\)

OpenStudy (solomonzelman):

yes, very good

OpenStudy (solomonzelman):

what are the remaining coordinates ?

OpenStudy (solomonzelman):

(you have 3 coordinates remaining)

OpenStudy (anonymous):

(4,-1)(5,2)(6,5)

OpenStudy (solomonzelman):

yes, very nice!

OpenStudy (solomonzelman):

so you will have (just for review) \(\large\color{black}{ \displaystyle (1,-10)}\) \(\large\color{black}{ \displaystyle (2,-7)}\) \(\large\color{black}{ \displaystyle (3,-4)}\) \(\large\color{black}{ \displaystyle (4,-1)}\) \(\large\color{black}{ \displaystyle (5,~~~2)}\) \(\large\color{black}{ \displaystyle (6,~~~5)}\)

OpenStudy (solomonzelman):

bye

OpenStudy (anonymous):

can i ask one more and you tell me if i did it right?

OpenStudy (anonymous):

Graph the six terms of a finite series where a1 = 5 and r = 1.25